Calculate the molal depression constant of a solvent, which has freezing point 16.6^0C and latent heat of fusion 180.75 J/g:
a]2.68      b]3.86

Grade:

1 Answers

Arun
4 years ago

Dependence of K_{f} -

K_{f}= \frac{RT_{f}^{2}m}{1ooo\Delta H_{fusion}}

-

 

 K_{f}=\frac{RT_{f}^{2}}{100\times L_{f}},\: R=8.314JK^{-1}mol^{-1}

T_{f}=273+16.6 = 289.6K;\: \: L_{f}=180.75Jg^{-1}

K_{f}=\frac{8.314 \times 289.6 \times 289.6}{1000\times 180.75}

 

On solving Kf = 3.86

Hope it helps

In case of any difficulty, Please feel free to ask again